OpenStudy (anonymous):

Find the exact real value of arcsin(1/2)

OpenStudy (anonymous):

help me please!

OpenStudy (anonymous):

arcsin(1/2)=30

OpenStudy (anonymous):

what about in pi form??@ASAAD123

OpenStudy (anonymous):

@ASAAD123

OpenStudy (anonymous):

(1/6)pi

jimthompson5910 (jim_thompson5910):

30 degrees = 30*(pi/180) radians 30 degrees = 30pi/180 radians 30 degrees = pi/6 radians

jimthompson5910 (jim_thompson5910):

to convert from degrees to radians, you multiply by pi/180

OpenStudy (anonymous):

thanks youu